OpenStudy (anonymous):

Determine the solutions to the quadratic equation -4x2 - 32x - 64 = 0.

OpenStudy (goformit100):

@nphuongsun93

OpenStudy (anonymous):

Help me @nphuongsun93

OpenStudy (anonymous):

@dpaInc Helllpp(:

OpenStudy (anonymous):

LOL @goformit100 -4x2 - 32x - 64 = 0 x^2 + 8x + 16 = 0 (x+4)^2 =0 x=-4

OpenStudy (goformit100):

he is coorect thanks @nphuongsun93